# LLM4Decompile: Decompiling Binary Code with Large Language Models

### v2

@@ 4.2.1 Experimental Setups
The training data is constructed using ExeBench, with Ghidra Headless employed to decompile the binary object file. Due to constraints in computational resources, only 400K functions each with optimization levels from O0 to O3 (1.6M samples, 1B tokens) are used for training and the evaluation is conducted on HumanEval-Decompile. The models are trained using the same template described in Section 4.1.1. In addition, following previous work Hosseini and Dolan-Gavitt (2022); Armengol-Estapé et al. (2023), we access the readability of decompiled results in terms of Edit Similarity score.
@@
Configuration settings for the model are consistent with those in Section 4.1.1. For the 1.3B, 6.7B models, the fine-tuning process involves 2B tokens in 2 epochs, and requires 2, and 8 days respectively on $1008\times A$ 100 respectively. Limited by the resource, for 33B model we only train for 200M tokens. For evaluation, we first access the re-executability rate of Ghidra to establish a baseline. Subsequently, GPT-4o is used to enhance Ghidra’s decompilation result with the prompt, Generate linux compilable C/C++ code of the main and other functions in the supplied snippet without using goto, fix any missing headers. Do not explain anything., following DecGPT Wong et al. (2023). Finally, we use LLM4Decompile-Ref models to refine the Ghidra’s output.
Model/Obfuscation Control Flow Flattening Bogus Control Flow O0 O1 O2 O3 AVG O0 O1 O2 O3 AVG LLM4Decompile-End-6.7B 0.0427 0.0488 0.0488 0.0305 0.0427 0.0976 0.0732 0.0793 0.0976 0.0869 Ghidra 0.1220 0.0671 0.0610 0.0671 0.0793 0.0610 0.0427 0.0305 0.0427 0.0442 +LLM4Decompile-Ref-6.7B 0.0671 0.0366 0.0488 0.0549 0.0519 0.1585 0.1402 0.0854 0.0793 0.1159
@@
4.2.2 Experimental Results
The results for the baselines and Refined-Decompile approaches are summarized in Table 4.1.2. For the pseudo-code decompiled by Ghidra, which is not optimized for re-execution, only an average of 20.1% of them pass the test cases. GPT-4o assists in refining this pseudo-code and enhancing its quality. The LLM4Decompile-Ref models offer substantial improvements over Ghidra’s outputs, with the 6.7B model yielding a 160% increase in re-executability. Similar to the discussion in Section 4.1.2, the 33B model outperforms the 1.3B model even though it used considerably less training data. And it achieves performance that is only 3.6% below the 6.7B model, which benefited from ten times more training data. When compared to LLM4Decompile-End-6.7B, the LLM4Decompile-Ref-6.7B model, though trained on just 10% of the data in LLM4Decompile-Ref models, shows a 16.2% performance increase, suggesting a greater potential for the Refined-Decompile approach.
An analysis of readability across different methods is also conducted and presented in Table 4.1.2, examples are presented in Figure 4. For text similarity, all decompiled outputs diverge from the original source code, with Edit Similarity ranging from 5.7% to 14.0%, primarily because the compilation process removes variable names and optimizes the logic structure. Ghidra generates pseudo-code that is particularly less readable with 6.2% Edit Similarity on average. Interestingly, with refinement from GPT (Ghidra+GPT-4o), there is a marginal decrease in Edit Similarity. GPT assists in refining type errors like undefined4 and ulong (Figure 4), however, it struggles to accurately reconstruct for loops and array indexing. In contrast, both LLM4Decompile-End and LLM4Decompile-Ref generate outputs that are more aligned with the format of the source code and easier to comprehend. To summarize, domain-specific fine-tuning is crucial for enhancing re-executability and readability of decompilation outputs.
@@ 5 Obfuscation Discussion
In software development, engineers typically implement obfuscation techniques before releasing binary files to the public. This is done to protect the software from unauthorized analysis or modification. In our study, we focus on two fundamental obfuscation techniques as suggested in Obfuscator-LLVM Junod et al. (2015): Control Flow Flattening (CFF) and Bogus Control Flow (BCF). These techniques are designed to disguise the true logic of the software, thereby making decompilation more challenging to protect the software’s intellectual property. We present the details of these two techniques in the Appendix C.
@@ Appendix A ExeBench Setups
For every sample in ExeBench’s executable splits, assembly code from *.s file—a compiler’s intermediate output as discussed in Section 3.1 and Figure 1—is required to compile the sample into a binary. The specific compilation settings and processing details, however, are not provided by the authors. Consequently, we choose to compile the code in a standard way and manage to compile only half of the samples. This leaves us with 443K out of 797K samples for the executable training set and 2621 out of 5000 samples for the executable test set. Accordingly, we train our model on the 443K samples and conduct the re-executability evaluation on these 2621 samples, the results are shown in Table 1.
The researchers from Slade Armengol-Estapé et al. (2023), who also developed ExeBench Armengol-Estapé et al. (2022), have published their decompilation findings on ExeBench. They chose to decompile the intermediate output, or assembly code from *.s file, directly without further compilation into binaries, where in practice, such intermediate output is rarely released by software developers. Their reported results, as seen in Table 5, show a significant difference from ours. Their version of ChatGPT achieved a re-executability rate of 22.2% and an edit similarity of 44.0% under O0 optimization. On the other hand, our GPT-4o model only reached a 4.4% re-executability rate and a 7.9% edit similarity. The approach taken by Slade involves settings not commonly available in practical decompilation scenarios, which explains why their results vary significantly from ours. We adheres to a more realistic setting, decompiling binary files based solely on their intrinsic data, without any external information.
@@
To further illustrate our settings, Figure 5 offers an example where the source function includes specific user-defined types like Ltc4151State, Ltc4151, and device. However, these types are completely lost after compilation, i.e., no information related to these user-definitions can be found in the binary (disassembled ASM code). Consequently, GPT-4o is unable to reconstruct these types based purely on the ASM (the realistic setting), instead converting them to default types int or pointer, producing non-executable code. This issue was pervasive across the ExeBench test set, leading to the failure of GPT-4o models in decompiling the ExeBench samples in a realistic setting.
@@
Appendix B Further Analysis of LLM4Decompile-Ref
@@
The error analysis presented in Figure 7 for LLM4Decompile-End-6.7B indicates that logical errors are prevalent in the HumanEval-Decompile scenarios, with 64% of errors due to assertions that the decompiled codes do not pass. In the ExeBench dataset, which features real functions with user-defined structures and types, the major challenges are related to reclaiming these user-specific components. Where 50% of the errors come from undeclared functions, and 28% from improper use of structures. Given that these user-defined details are typically lost during the compilation process, reconstructing them can be particularly challenging. Integrating techniques like Retrieval Augmented Generation might supplement the decompilation process with necessary external information.
Figure 6: Re-executability rate with the growth of input length. 6.7B model is more robust against input length.
Appendix C Obfuscation Techniques
@@
We provide the details of two classic obfuscation techniques suggested in Obfuscator-LLVM.
Control Flow Flattening
@@
enhances the security of software by transforming its straightforward, hierarchical control flow into a more complex, flattened structure. The workflow involves breaking a function into basic blocks, arranging these blocks at the same level, and encapsulating them within a switch statement inside a loop.
Bogus Control Flow
@@
modifies a function’s execution sequence by inserting an additional basic blockprior to the existing one. This added block includes an opaque predicate, followed by a conditional jump that leads back to the original block. Additionally, the original basic block is polluted with randomly selected, meaningless instructions.

### v3

@@ 4.2.1 Experimental Setups
The training data is constructed using ExeBench, with Ghidra Headless Ghidra (2024b) employed to decompile the binary object file. Due to constraints in computational resources, only 400K functions each with optimization levels from O0 to O3 (1.6M samples, 1B tokens) are used for training and the evaluation is conducted on HumanEval-Decompile. The models are trained using the same template described in Section 4.1.1. In addition, following previous work Hosseini and Dolan-Gavitt (2022); Armengol-Estapé et al. (2023), we access the readability of decompiled results in terms of Edit Similarity score.
@@
Configuration settings for the model are consistent with those in Section 4.1.1. For the 1.3B and 6.7B models, the fine-tuning process involves 2B tokens in 2 epochs, and requires 2 and 8 days respectively on $1008\times A$ 100. Limited by the resource, for 33B model we only train for 200M tokens. For evaluation, we first access the re-executability rate of Ghidra to establish a baseline. Subsequently, GPT-4o is used to enhance Ghidra’s decompilation result with the prompt, Generate linux compilable C/C++ code of the main and other functions in the supplied snippet without using goto, fix any missing headers. Do not explain anything., following DecGPT Wong et al. (2023). Finally, we use LLM4Decompile-Ref models to refine the Ghidra’s output.
4.2.2 Experimental Results
The results for the baselines and Refined-Decompile approaches are summarized in Table 4.1.2. For the pseudo-code decompiled by Ghidra, which is not optimized for re-execution, only an average of 20.1% of them pass the test cases. GPT-4o assists in refining this pseudo-code and enhancing its quality. The LLM4Decompile-Ref models offer substantial improvements over Ghidra’s outputs, with the 6.7B model yielding a 160% increase in re-executability. Similar to the discussion in Section 4.1.2, the 33B model outperforms the 1.3B model even though it used considerably less training data. And it achieves performance that is only 3.6% below the 6.7B model, which benefited from ten times more training data. When compared to LLM4Decompile-End-6.7B, the LLM4Decompile-Ref-6.7B model, though trained on just 10% of the data in LLM4Decompile-Ref models, shows a 16.2% performance increase, suggesting a greater potential for the Refined-Decompile approach. We present further analysis in Appendix D.
An analysis of readability across different methods is also conducted and presented in Table 4.1.2 with illustrative examples presented in Figure 4. For text similarity, all decompiled outputs diverge from the original source code, with Edit Similarity ranging from 5.7% to 14.0%, primarily because the compilation process removes variable names and optimizes the logic structure. Ghidra generates pseudo-code that is particularly less readable with 6.2% Edit Similarity on average. Interestingly, with refinement from GPT (Ghidra+GPT-4o), there is a marginal decrease in Edit Similarity. GPT assists in refining type errors like undefined4 and ulong (Figure 4). However, it struggles to accurately reconstruct for loops and array indexing. In contrast, both LLM4Decompile-End and LLM4Decompile-Ref generate outputs that are more aligned with the format of the source code and easier to comprehend. To summarize, domain-specific fine-tuning is crucial for enhancing re-executability and readability of decompilation outputs.
We further employed GPT-4o to evaluate readability Wang et al. (2023); Liu et al. (2023). Specifically, we guide GPT to assess syntax similarity (variables, loops, conditions) and structural integrity (logic flow, structure) using a structured template. We then summarize readability with a score from 1 (Poor) to 5 (Excellent), based on detailed comparisons between original and decompiled code. The template is available on our GitHub repository [^3] https://github.com/albertan017/LLM4Decompile/blob/main/samples/readability_template.txt. Table 4 summarizes our readability assessments on HumanEval-Decompile across various models and optimization levels.
Optimization Level O0 O1 O2 O3 AVG GPT-4o 2.8171 2.3537 2.2927 2.311 2.4436 Ghidra 2.9756 2.4085 2.5183 2.3841 2.5716 LLM4Decompile-End-6.7B 4.0732 3.4634 3.4024 3.2378 3.5442
Table 4: Evaluation by GPT-4o on the readability of decompiled results from various methods.
Compared with the results in Table 4.1.2, it indicates that Edit Similarity (ES) follows a trend similar to GPT evaluation. Although ES is mathematically based, its values can be difficult to interpret. For instance, a 15 ES score obtained by LLM4Decompile model may seem low, yet the decompiled function and the source code are highly aligned. In contrast, GPT evaluation, which measures readability conceptually, is more intuitive. A score of 4 on the GPT scale suggests that the decompiled code is nearly identical to the original. Nonetheless, these scores are derived from GPT’s "subjective" judgments. Combining insights from both ES and GPT-Eval could lead to a more thorough assessment of code readability.
Model/Obfuscation Control Flow Flattening Bogus Control Flow O0 O1 O2 O3 AVG O0 O1 O2 O3 AVG LLM4Decompile-End-6.7B 4.27 4.88 4.88 3.05 4.27 9.76 7.32 7.93 9.76 8.69 Ghidra 12.20 6.71 6.10 6.71 7.93 6.10 4.27 3.05 4.27 4.42 +LLM4Decompile-Ref-6.7B 6.71 3.66 4.88 5.49 5.19 15.85 14.02 8.54 7.93 11.59
@@ 5 Obfuscation Discussion
In software development, engineers typically implement obfuscation techniques before releasing binary files to the public Lachaux et al. (2021); Junod et al. (2015). This is done to protect the software from unauthorized analysis or modification. In our study, we focus on two fundamental obfuscation techniques as suggested in Obfuscator-LLVM Junod et al. (2015): Control Flow Flattening (CFF) and Bogus Control Flow (BCF). These techniques are designed to disguise the true logic of the software, thereby making decompilation more challenging to protect the software’s intellectual property. We present the details of these two techniques in the Appendix E.
@@ 6 Conclusions
Acknowledgments
This work is partially supported by the National Natural Science Foundation of China (No. 62372220), the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU/25200821), the NSFC Young Scientists Fund (Project No. 62006203), the Innovation and Technology Fund (Project No. PRP/047/22FX), and PolyU Internal Fund from RC-DSAI (Project No. 1-CE1E).
@@ Appendix A ExeBench Setups
For every sample in ExeBench’s executable splits, assembly code from *.s file—a compiler’s intermediate output as discussed in Section 3.1 and Figure 1—is required to compile the sample into a binary. The specific compilation settings and processing details, however, are not provided by their authors. Consequently, we choose to compile the code in a standard way and manage to compile only half of the samples. This leaves us with 443K out of 797K samples for the executable training set and 2621 out of 5000 samples for the executable test set. Accordingly, we train our model on the 443K samples and conduct the re-executability evaluation on these 2621 samples, the results are shown in Table 1.
The researchers from Slade Armengol-Estapé et al. (2023), who also developed ExeBench Armengol-Estapé et al. (2022), have published their decompilation findings on ExeBench. They chose to decompile the intermediate output, or assembly code from *.s file, directly without further compilation into binaries, where in practice, such intermediate output is rarely released by software developers. Their reported results, as seen in Table 6, show a significant difference from ours. Their version of ChatGPT achieved a re-executability rate of 22.2% and an edit similarity of 44.0% under O0 optimization. On the other hand, our GPT-4o model only reached a 4.4% re-executability rate and a 7.9% edit similarity. The approach taken by Slade involves settings not commonly available in practical decompilation scenarios, which explains why their results vary significantly from ours. We adhere to a more realistic setting, decompiling binary files based solely on their intrinsic data, without any external information.
@@
To further illustrate our settings, Figure 5 offers an example where the source function includes specific user-defined types like Ltc4151State, Ltc4151, and device. However, these types are completely lost after compilation, i.e., no information related to these user-definitions can be found in the binary (disassembled ASM code). Consequently, GPT-4o is unable to reconstruct these types based purely on the ASM (the realistic setting), but converting them to default types int or pointer, producing non-executable code. This issue was pervasive across the ExeBench test set, leading to the failure of GPT-4o models in decompiling the ExeBench samples in a realistic setting.
@@
Appendix B Compilable and Executable Binary
The statistics of training and testing datasets is summarized in the Table 7. We also present an example to illustrate the difference between these two datasets.
Dataset/Code ASM SRC Train-Executable 205.08 119.35 Test-Exebench 280.27 162.68 Train-Compilable 711.89 241.16 Test-Decompile-Eval 808.07 186.84
Table 7: Statistics of training and testing set.
Figure 6: Compilable data and Executable data.
As shown in Figure 6, the primary distinction between a compilable binary and an executable binary is the handling of function operation addresses. In a compilable file, the address for a jump operation is placeholder, representing only a relative offset within the function. Conversely, in an executable file, this jump operation address is resolved during the linking process, pointing directly to the specific memory location where the code will execute.
Appendix C Further Analysis of LLM4Decompile-End
Figure 7: Re-executability rate with the growth of input length. The 6.7B model is more robust against input length.
@@ Appendix C Further Analysis of LLM4Decompile-End
The error analysis presented in Figure 8 for LLM4Decompile-End-6.7B indicates that logical errors are prevalent in the HumanEval-Decompile scenarios, with 64% of errors due to assertions that the decompiled codes do not pass. In the ExeBench dataset, which features real functions with user-defined structures and types, the major challenges are related to reclaiming these user-specific components. Where 50% of the errors come from undeclared functions, and 28% from improper use of structures. Given that these user-defined details are typically lost during the compilation process, reconstructing them can be particularly challenging. Integrating techniques like Retrieval Augmented Generation Lewis et al. (2020b) might supplement the decompilation process with necessary external information.
Appendix D Data Quality, Volume and Model
Data Quality
In this project, we intentionally limited our data preprocessing to classical techniques such as filtering short texts and removing duplicates. This approach was chosen to establish a fair baseline model for decompilation that minimizes potential biases, aiming to provide a broad, unrefined baseline model that reflects diverse scenarios. We acknowledge that selective data removal, specifically excluding data incompatible with standard C libraries Free Software Foundation (2024), can enhance performance, as evidenced in Table 8 with Decompile-Eval, which only relies on standard C libraries. While refining the dataset can lead to improved performance, our primary goal in this study was to set a foundational baseline for the community. We believe this baseline can serve as a starting point encouraging future research to refine and expand.
Models Re-executability Rate Optimization Level O0 O1 O2 O3 AVG Compilable-6.7B 51.83 33.54 32.32 32.32 37.50 +Executable (2B tokens) 68.05 39.51 36.71 37.20 45.37 +Exe w. Standard C (100M tokens) 71.80 42.68 41.31 41.46 49.31
Table 8: Performance improves when training data (+Exe w. Standard C) closely resembles the testing set patterns.
Data Volume
We have summarized the relationship between performance and training epochs (2B token in one epoch) for the LLM4Decompile-Ref-1.3B model in Table 9. From this, it is clear that a single epoch serves as a strong baseline, while two epochs optimize performance. Additional epochs tend to lead to overfitting and diminished results.
Epoch O0 O1 O2 O3 Avg 1 67.68 41.46 41.46 35.37 46.49 2 68.29 40.85 40.85 37.20 46.80 3 62.80 37.80 36.59 29.88 41.77 4 51.22 32.93 27.44 26.22 34.45
Table 9: Performance of LLM4Decompile-Ref-1.3B on HumanEval-Decompile w.r.t. different training epochs.
Moreover, we have included results from the 6.7B model to address scaling issues and provided comparison with the 1.3B and 33B models in Table 10. Notably, the 6.7B model achieves comparable performance to the 1.3B model with only 20% of the data, and the 33B model reaches similar outcomes to the 6.7B with just 10% of the data, although these ratios may differ with varying datasets.
Size epoch O0 O1 O2 O3 Avg 1.3B 1.0 67.68 41.46 41.46 35.37 46.49 6.7B 0.05 55.49 31.71 34.76 30.49 38.11 6.7B 0.1 57.93 36.74 32.77 32.01 39.86 6.7B 0.2 65.85 37.80 40.24 34.76 44.66 6.7B 0.5 65.55 45.73 43.29 43.75 49.58 6.7B 1.0 72.56 45.73 43.90 42.68 51.22 33B 0.1 70.73 47.56 43.90 41.46 50.91
Table 10: Performance of LLM4Decompile-Ref models on HumanEval-Decompile w.r.t. different training epochs.
Our findings suggest that model scaling can significantly enhance performance when the training data is adequately large (100M tokens), but repetitive training risks overfitting after a few epochs.
Model
Choosing the right base model for decompilation training significantly influences performance. Our first choice, Deepseek-Coder-6.7B, delivered an encouraging average re-executability rate of 52.74% on the HumanEval-Decompile benchmark. Conversely, LLM-Compiler-7B Cummins et al. (2024), trained to compile source code into LLVM IR Lattner and Adve (2004)—the opposite of decompilation—served as a more effective foundation, enhancing performance by 3.5% compared to Deepseek-Coder-6.7B. Additionally, Yi-Coder-9B 01-AI (2024), introduced in September 2024 as the current state-of-the-art model, markedly improved decompilation training results by 23.1%. Furthermore, CodeStral-22B Mistral-AI (2024), benefiting from its larger architecture, provided a 21.7% improvement over smaller models.
Model/Opt. Level O0 O1 O2 O3 AVG DeepSeek-Coder-6.7B 74.39 46.95 47.56 42.07 52.74 LLM-Compiler-7B 72.56 51.83 48.78 45.12 54.57 Yi-Coder-9B 79.27 62.20 61.59 56.71 64.94 CodeStral-22B 80.49 58.54 59.76 57.93 64.18
Table 11: Comparison of re-executability rates for base models in the LLM4Decompile-Ref series.
Appendix E Obfuscation Techniques
@@
We provide the details of two classic obfuscation techniques suggested in Obfuscator-LLVM Junod et al. (2015).
Control Flow Flattening.
@@
It enhances the security of software by transforming its straightforward, hierarchical control flow into a more complex, flattened structure. The workflow involves breaking a function into basic blocks, arranging these blocks at the same level, and encapsulating them within a switch statement inside a loop.
Bogus Control Flow.
@@
It modifies a function’s execution sequence by inserting an additional basic block prior to the existing one. This added block includes an opaque predicate, followed by a conditional jump that leads back to the original block. Additionally, the original basic block is polluted with randomly selected, meaningless instructions.